find the perimeter of a triangular garden that measures 10ft x 10ft by (2x+3)feet. Use P =a+b+c
1
Expert's answer
2014-09-12T12:47:59-0400
If lengths of sides are a=10ft, b=10ft, c=(2x+3)feet, then perimeter is P=a+b+c=10+10+2x+3=2x+23 feet